Artificial satellite revolves around the earth in 2.5 hours in a circular orbit . Find the height of satellite above the earth assuming earth as a sphere of radius 6370 kilometre.
1
Expert's answer
2019-04-26T11:45:06-0400

The Kepler's third law states

"T^2=\frac{4\pi^2 a^3}{GM}"

So, the radius of satellite orbit

"a=\sqrt[3]{\frac{GMT^2}{4\pi^2}}"

"=\sqrt[3]{\frac{6.67\times 10^{-11}\times 5.97\times 10^{24}\times(2.5\times 3600)^2}{4\pi^2}}"

"=9.352\times 10^6\:\rm{m}"

The height of satellite above the earth surface

"H=a-R"

"=9.352\times 10^6-6.370\times 10^6=2.982\times 10^6\:\rm{m}"

"=2982\:\rm{km}"